### v1

@@ 5.1 LLVM Phase Ordering POMDP
We formulate the phase ordering problem of LLVM as a partially observable Markov decision process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},r,p,\mu,\mathcal{O},\phi)$. The state space $\mathcal{S}$ covers all possible Intermediate Representations (IRs), the action space $\mathcal{A}$ comprises individual compiler optimization passes, and the reward function $r$ is defined by the metric being optimized. The transition dynamics $p:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ represents the outcome of applied IR transformations. The initial state distribution $\mu\in\Delta(\mathcal{S})$ captures all IRs of interest, which can be approximated via uniform sampling from the training dataset. The observation function $\phi:\mathcal{S}\mapsto\mathcal{O}$ maps the underlying IR into the observation space. Complex observation spaces offer comprehensive program information, while expert-designed features embed problem-specific knowledge, potentially enhancing the optimization policy’s generalization by eliminating irrelevant details. The code optimization agent with a policy $\pi$ interacts with the compiler environment according to the protocol described in Section 2.1. We fix the horizon of interactions to 45 steps.
Following Autophase huang2019autophase and CompilerGym cummins2022compilergym , the action space consists of 42 optimization passes out of over a hundred in LLVM, which effectively prunes the vast optimization space while not sacrificing the performance of the learned policy. The observation space is a concatenation of two numeric features: the 56-dimension Autophase feature summarizing the statistics of the IR, such as the number of specific basic blocks, branches, and instructions, and a 42-dimension histogram of the agent’s previous actions. The reward function is defined as the normalized change of the optimization metric $C(s)$:
$r_{t+1}=\frac{C(s_{t})-C(s_{t+1})}{C(s_{0})-C(s_{\mathrm{b}})},$ (2)
where a lower $C$ indicates better performance and $C(s_{\mathrm{b}})$ stands for the performance of a baseline policy, such as built-in -Oz or -O3. A total reward greater than 1 means that the optimization sequence performs better than the baseline policy.
5.2 Training Dataset
We construct our training datasets, by considering three datasets from distinct domains. The CodeContests dataset li2022competition consists of over 13,000 competitive problems, each on average having hundreds of solutions in three languages. We subsample up to ten C++ solutions for each training problem, resulting in 110,240 programs, as part of our training data, and sample one solution for each of 100 test problems from CodeContests as our validation data. We also include the full Form-AI dataset tihanyi2023formai with 112,000 AI-generated programs and the collection of the largest 9,998 single-function programs and 15,264 multiple-function programs from the AnghaBench da2021anghabench into our training data.
5.3 Model-based RL Method
We utilize a deep model-based RL method, DreamerV3 hafner2023mastering , to model the compiler environment upon which it learns an effective policy that maps observations to actions with parameterized deep neural networks.
World Model
@@ 5.3 Model-based RL Method
The world model simulating the compiler environment is formulated as a latent dynamics model with the following four components:
$&\text{Representation model:}&&z_{t}\sim q_{\theta}(z_{t}\mid z_{t-1},a_{t-1},o_{t})\\ &\text{Transition model:}&&\hat{z}_{t}\sim p_{\theta}(\hat{z}_{t}\mid z_{t-1},a_{t-1})\\ &\text{Image decoder:}&&\hat{o}_{t}\sim p_{\theta}(\hat{o}_{t}\mid z_{t})\\ &\text{Reward decoder:}&&\hat{r}_{t}\sim p_{\theta}(\hat{r}_{t}\mid z_{t})\end{aligned}$ (3)
The representation model estimates the latent state $z_{t}$ based on the previous state $z_{t-1}$, the previous action $a_{t-1}$ and the current observation $o_{t}$, while the transition model predicts it directly from $z_{t-1}$ and $a_{t-1}$. The overall models are jointly learned by minimizing the negative evidence lower bound (ELBO) jordan1999introduction ; kingma2013auto :
$\mathcal{L}_{\text{model}}(\theta)\doteq\;\mathbb{E}_{q_{\theta}\left(z_{1:T}\mid a_{1:T},o_{1:T}\right)}$ $\Big{[}\sum_{t=1}^{T}\Big{(}{-\ln p_{\theta}(o_{t}\mid z_{t})-\ln p_{\theta}(r_{t}\mid z_{t})}$ (4)
$\,\,\,\,{+\,\text{KL}\left[q_{\theta}(z_{t}\mid z_{t-1},a_{t-1},o_{t})\,\|\,p_{\theta}(\hat{z}_{t}\mid z_{t-1},a_{t-1})\right]}\Big{)}\Big{]}.$
@@
In practice, we employ a reward smoothing technique lee2023dreamsmooth to mitigate the sparsity and the long tail distribution of rewards during an episode. This is achieved through the following equation:
@@
Actor-Critic Learning
The actor and critic neural networks are parameterized on top of the latent representations:
@@
The actor and critic are jointly trained on the same imagined trajectories $\{\hat{z}_{\tau},\hat{a}_{\tau},\hat{r}_{\tau}\}$ with horizon $H$, generated by the interactions between the transition model and reward model in Eq. (3) and the actor: starting at the latent state $\hat{z}_{t}=z_{t}$, at each step $\tau=t,t+1,t+2,\dots$, the policy takes an action $\hat{a}_{\tau}\sim\pi_{\psi}\left(\hat{a}_{\tau}\mid\hat{z}_{\tau}\right)$, and transits to the next latent state $\hat{z}_{\tau+1}\sim p_{\theta}(\hat{z}_{\tau+1}\mid z_{\tau},a_{\tau})$ with a reward $\hat{r}_{\tau+1}\sim p_{\theta}(\hat{r}_{\tau+1}\mid\hat{z}_{\tau+1})$. The critic is trained to predict the $\lambda$-return sutton2018reinforcement through a discrete regression loss bellemare2017distributional ; hafner2023mastering :
@@
$V_{\tau}^{\lambda}\doteq\hat{r}_{\tau}+\gamma\begin{cases}(1-\lambda)v_{\xi}(\hat{z}_{\tau+1})+\lambda V_{\tau+1}^{\lambda}&\text{if}\ \tau<t+H\\ v_{\xi}(\hat{z}_{\tau+1})&\text{if}\ \tau=t+H.\end{cases}$ (8)
The actor, meanwhile, is trained to maximize the imagined return through the REINFORCE policy gradient williams1992simple :
@@
$\mathcal{L}_{\text{actor}}(\psi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\left[\sum^{t+H}_{\tau=t}\left(-\left(V_{\tau}^{\lambda}-v_{\xi}\left(\hat{z}_{\tau}\right)\right)\log\pi_{\psi}\left(\hat{a}_{\tau}\mid\hat{z}_{\tau}\right)-\eta\,\text{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]\right)\right],$ (9)
@@
where $\text{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]$ is an entropy regularization which encourages exploration, and $\eta$ is a hyperparameter that adjusts the regularization strength.
Appendix A Related Work
In this section, we will elaborate a more extensive review of the development of machine learning methods for code optimization in compilers.
One of the key challenges for compilation is to determine which code transformations to apply, how to apply them (e.g., using suitable parameters), and in what order. This requires effectively searching and evaluating a massive number of possible options, known as iterative compilation bodin1998iterative or autotuning datta2008stencil . However, this search-based approach only finds a good optimization for one specific program and does not generalize into a compiler heuristic. This limitation underscores the importance of integrating machine learning techniques.
Pioneering work has delved into supervised machine learning, adopting two main approaches. The first one learns a predictive model that can directly predict the best option. It involves iteratively compiling training programs to identify the most effective compilation strategy for each, which then serves as the labels of training data. An early example calder1997evidence used a neural network for branch prediction, and one more well-known work is MilepostGCC fursin2008milepost , a practical attempt to integrate machine learning into a production compiler, GCC. It employs models trained on a large dataset of programs distributed over the Internet. The second approach aims to learn a cost or performance function capable of estimating the quality of various compiler options, which enables evaluation of a range of possible options without the need to compile and profile each one stephenson2003meta ; luk2009qilin .
Recent advancements have seen reinforcement learning (RL) techniques making strides in compiler optimization, circumventing the requirement for collecting optimal labeled data kulkarni2012mitigating . This technique has been applied to optimize individual compilation heuristics, such as inlining trofin2021mlgo , loop transformation haj2020neurovectorizer ; brauckmann2021reinforcement , and graph partitioning mirhoseini2017device . Several works relevant to us have explored the full optimization pipeline, i.e. the LLVM phase ordering problem, including AutoPhase huang2019autophase , CORL mammadli2020static , and POSET-RL jain2022poset . These methods predominantly utilize model-free RL algorithms while our work pioneers the use of an advanced model-based RL approach to reduce real compiler interactions. Model-based RL sutton1991dyna learns a simulation model of the compiler environment, which is related in spirit to the aforementioned approach of supervised performance models. However, it goes a step further by also learning a policy capable of directly determining the best optimization option, thereby eliminating the necessity for a guided search process.

### v2

@@ 3.1 Phase Ordering Decision Process
As illustrated in Figure 1, one key problem of compiler optimization is to find the optimal sequence of optimization passes for a given program, which is also referred to as the phase ordering problem. It can be naturally formulated as a partially observable Markov decision process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},r,p,\mu,\mathcal{O},\phi)$ (Sutton and Barto 2018). The state space $\mathcal{S}$ covers all possible Intermediate Representations (IRs), the action space $\mathcal{A}$ comprises individual compiler optimization passes, and the reward function $r$ is defined by the metric being optimized. The transition dynamics $p:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ represents the outcome of applied IR transformations. The initial state distribution $\mu\in\Delta(\mathcal{S})$ captures all IRs of interest, which can be approximated via uniform sampling from the training dataset. The observation function $\phi:\mathcal{S}\mapsto\mathcal{O}$ maps the underlying IR into the observation space. Complex observation spaces offer comprehensive program information, while expert-designed features (Haj-Ali et al. 2020) embed problem-specific knowledge, potentially enhancing the optimization policy’s generalization by eliminating irrelevant details.
A code optimization agent guided by policy $\pi$ determines the optimization sequence of an input program through a series of interactions with the compiler environment. At each time step $t=0,1,2,\dots$, the agent applies an action $a_{t}$ to current IR based on observation $o_{t}$ and its policy $\pi(a_{t}\mid o_{\leq t})$, which transforms the current state $s_{t}$ to the next state $s_{t+1}=p(s_{t},a_{t})$ and will receive a reward $r_{t}$. The goal of the agent is to learn an optimization policy $\pi$ that effectively maximizes the cumulative rewards $G=\sum_{t}r_{t}$, thereby optimizing the performance metrics of the final IR.
3.2 Learning a Compiler World Model
We develop our CompilerDream approach by adapting an advanced model-based RL method, Dreamer (Hafner et al. 2023), as depicted in Figure 2. This approach involves learning a predictive world model $(\hat{p}_{\theta},\hat{r}_{\theta})$ of the compiler environment parameterized by $\theta$ that approximates the underlying transition dynamics $p(o_{t+1}|o_{\leq t},a_{\leq t})$ and reward function $r(o_{\leq t},a_{\leq t})$.
@@ 3.2 Learning a Compiler World Model
The compiler world model simulating the compiler environment is formulated with the following four components:
Representation model: $z_{t}\sim q_{\theta}(z_{t}\mid z_{t-1},a_{t-1},o_{t})$ (1)
Transition model: $\hat{z}_{t}\sim p_{\theta}(\hat{z}_{t}\mid z_{t-1},a_{t-1})$
Observation decoder: $\hat{o}_{t}\sim p_{\theta}(\hat{o}_{t}\mid z_{t})$
Reward decoder: $\hat{r}_{t}\sim p_{\theta}(\hat{r}_{t}\mid z_{t}).$
The representation model estimates a neural compiler state $z_{t}$ from the current observation $o_{t}$ of real compiler state, the previous state $z_{t-1}$ and the previous optimization action $a_{t-1}$. A representation loss $\mathcal{L}_{\mathrm{repr}}$ is used to train the neural compiler states to accurately reconstruct the observation and reward by two decoders:
$\mathcal{L}_{\mathrm{repr}}(\theta)\doteq-\ln p_{\theta}(o_{t}|z_{t})-\ln p_{\theta}(r_{t}|z_{t}).$ (2)
The transition model captures the dynamics of the compiler world model, predicting future neural state ${\hat{z}}_{t}$ directly from $z_{t-1}$ and $a_{t-1}$. A prediction loss minimizes the difference between the estimated state $z_{t}$ and the predicted state ${\hat{z}}_{t}$, simultaneously enhancing the transition model’s accuracy in predicting future states and making the state produced by the representation model easier to predict:
$\mathcal{L}_{\mathrm{pred}}(\theta)\doteq\mathrm{KL}\left[q_{\theta}(z_{t}|z_{t-1},a_{t-1},o_{t})\,\|\,p_{\theta}(\hat{z}_{t}|z_{t-1},a_{t-1})\right].$ (3)
The overall models are jointly learned by minimizing the sum of the representation loss and the prediction loss.
Compiler simulation
We can simulate the behavior of the real compiler using our trained compiler world model. Specifically, a simulated compiler optimization trajectory $\{\hat{z}_{\tau},\hat{a}_{\tau},\hat{r}_{\tau},\hat{o}_{\tau}\}$ with horizon $H$ can be generated by the interactions between the world model and an optimization agent: starting at a neural state $\hat{z}_{t}\sim q_{\theta}(z_{t}|o_{t})$, at each step $\tau=t,t+1,t+2,\dots$, the agent takes an action $\hat{a}_{\tau}\sim\pi_{\psi}\left(\hat{a}_{\tau}|\hat{z}_{\tau}\right)$, and transits to the next latent state $\hat{z}_{\tau+1}\sim p_{\theta}(\hat{z}_{\tau+1}|z_{\tau},a_{\tau})$ with a reward $\hat{r}_{\tau+1}\sim p_{\theta}(\hat{r}_{\tau+1}|\hat{z}_{\tau+1})$. Predicted real compiler state $\hat{o}_{\tau+1}$ can be optionally reconstructed by the observation decoder.
Reward smoothing
@@
During an optimization sequence, not all passes can effectively transform the code, and the improvements gradually saturate. We employ a reward smoothing technique (Lee, Abbeel, and Lee 2024) to mitigate the sparsity and the long tail distribution of rewards during an episode. This is achieved through the following:
@@
3.3 Learning a Code Optimization Agent
We then train a code optimization agent on top of the world model, which consists of an actor and a critic neural networks parameterized on top of the neural compiler state:
@@ 3.3 Learning a Code Optimization Agent
The critic evaluates the cumulative optimization return that the actor can achieve starting from a given neural state, also known as value or return, denoted as $v_{\xi}(\hat{z}_{t})$. This is trained by minimizing the difference between the predicted value $v_{\xi}(\hat{z}_{t})$ and the Monte-Carlo return or more advanced bootstrapped $\lambda$-return (Sutton and Barto 2018):
@@
$V_{\tau}^{\lambda}\doteq\hat{r}_{\tau}+\gamma\begin{cases}(1-\lambda)v_{\xi}(\hat{z}_{\tau+1})+\lambda V_{\tau+1}^{\lambda}&\text{if}\ \tau<t+H\\ v_{\xi}(\hat{z}_{\tau+1})&\text{if}\ \tau=t+H.\end{cases}$
The actor produces a policy $\pi_{\psi}$ that predicts the best optimization pass to choose, represented as an action distribution, which is trained to maximize the simulated return through the REINFORCE policy gradient (Williams 1992):
@@
$(\displaystyle\mathcal{L}_{\text{actor}}(\psi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\Big{[}\sum^{t+H}_{\tau=t}\Big{(}$$-\left(V_{\tau}^{\lambda}-v_{\xi}(\hat{z}_{\tau})\right)\log\pi_{\psi}(\hat{a}_{\tau}|\hat{z}_{\tau})$
$-\eta\,\mathbb{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]\Big{)}\Big{]},$ (7)
@@
where $\mathbb{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]$ is an entropy regularization which encourages exploration, and $\eta$ is a hyperparameter that adjusts the regularization strength.
Both the actor and critic are completely trained on world model-simulated compiler optimization trajectories.
3.4 Data Curation
To facilitate that our CompilerDream method can effectively generalize to unseen situations, a concept known as zero-shot generalization, we have identified three critical factors in preparing our training dataset. First, the dataset should reflect naturalness. During our preliminary experiments, we found that those generated code datasets like Csmith (Yang et al. 2011) and llvm-stress (Lattner and Adve 2004) provide no benefits or even hurt the generalization to real-world scenarios. Second, the dataset should be large to prevent the agent from overfitting to a small number of codes and failing to generalize. Last, the code must exhibit high quality from an optimization perspective. We need data with complex algorithmic logic and potential for optimizations to help the world model and the agent better understand the problem’s intricacies. Datasets like AnghaBench (Da Silva et al. 2021) which consist of millions of human-written codes collected from GitHub, are often too simple to allow for significant optimization improvements.
Therefore, we choose to construct our training datasets on top of the CodeContests dataset released with AlphaCode (Li et al. 2022), which consists of over 13,000 problems of coding competition, and each problem, on average, has hundreds of solutions in multiple languages. We subsample up to ten C++ solutions for each training problem, resulting in 110,240 programs, as our training data, and sample one solution for each of 100 test problems as our validation data.
4 Experiments
We evaluate CompilerDream in three different settings and compare its performance with prior state-of-the-art methods. We first demonstrate that, without considering generalization, CompilerDream excels as a powerful autotuning method, leading the CompilerGym leaderboard (Section 4.2). Then we show that CompilerDream can learn a general compiler world model capable of accurately estimating and selecting pass sequences (Section 4.3) and finally learn a general optimization agent that directly generates a pass sequence tailored to any specific input program. (Section 4.4)
4.1 Evaluation
Our experiments focus on code size reduction, which benefits applications targeting low-resource hardware such as embedded systems. This focus is driven by the practical advantages of code size as a metric: it is cost-effective to construct extensive compilable training and test datasets and to evaluate the optimization performance for code size.
Metric
To be more robust to outliners, we evaluate the code size optimization results by the geometric mean of IR instruction count reduction of program $s$ in a dataset $\mathcal{D}$:
$R_{\mathcal{D}}(\pi)=\left(\prod_{s\in\mathcal{D}}\frac{C(s_{Oz})}{C(s_{t}^{\pi})}\right)^{\frac{1}{|\mathcal{D}|}}$ (8)
where $C$ counts the number of IR-instructions, $s_{Oz}$ is the IR optimized by LLVM -Oz flag, and $s_{t}^{\pi}$ is the final IR produced by the algorithm $\pi$. A value above $1$ indicates that the algorithm performs better than -Oz option on average.
Benchmarks
We evaluate our method mainly on benchmarks from the CompilerGym platform (Cummins et al. 2022): benchmark suites including cBench (Fursin et al. 2007), CHStone (Hara et al. 2008), MiBench (Guthaus et al. 2001), and NASA Parallel Benchmarks (NPB) (Bailey et al. 1995), in addition to kernels from open source projects including BLAS (Lawson et al. 1979), Linux, OpenCV (Culjak et al. 2012), and TensorFlow (Abadi et al. 2016). We disregard benchmarks generated by synthetic program generators (Lattner and Adve 2004; Yang et al. 2011) since they do not align with real-world scenarios. We follow the standard data splits of CompilerGym.
Additionally, we evaluate CompilerDream on a large-scale dataset named FormAI (Tihanyi et al. 2023). FormAI comprises a vast collection of AI-generated C programs with diverse functionalities and coding styles. We filtered out codes that failed to compile into CompilerGym benchmarks, resulting in a test set of 109,016 programs.
Inspired by FormAI, we also constructed a dataset of 50 Objective-C programs generated by a Large Language Model to further evaluate CompilerDream’s ability to generalize to different programming languages. All details on benchmarks can be found in Appendix C.
Figure 3: Generel code optimization results: Code size reduction in terms of IR instruction count over LLVM -Oz under different methods. Bars indicate the geometric mean and min-max range across test programs in each benchmark dataset.
4.2 Autotuning: CompilerGym Leaderboard
Method Walltime
Code size
Reduction
CompilerDream
   + Guided Search
60.8s 1.073$\times$
PPO + Guided Search 69.8s 1.070$\times$
CompilerDream 2.9s 1.068$\times$
Random Search ($t$=10800) 10,512.4s 1.062$\times$
Random Search ($t$=3600) 3,630.8s 1.061$\times$
Greedy Search 169.2s 1.055$\times$
GATv2 + DD-PPO 258.1s 1.047$\times$
Table 1: Autotuning results: CompilerDream’s code size reduction and time consumption on cBench dataset compared with other methods from CompilerGym leaderboard (Facebook 2022). Results for methods other than ours are taken directly from the leaderboard, where learning-based methods report only inference time as wall time.
We first use CompilerDream as an autotuning method to verify that its training process has sufficient capability to discover high-quality optimization sequences, supporting our goal of training an effective general code optimization agent.
Implementation
We target the task of CompilerGym leaderboard (Facebook 2022), optimizing pass sequences across 23 cBench programs. CompilerDream is trained for 25 hours, averaging about one hour per program—a time budget comparable to search methods. After training, besides evaluating single-trial optimization performance (more in Section 4.4), we test CompilerDream+Guided Search, following the leaderboard’s top method. This approach uses the agent’s policy $\pi_{\psi}(a_{t}|s_{t})$ to guide random search and limits search time to 1 minute per program (details in Appendix B.7).
Results
As shown in Table 1, CompilerDream’s policy achieves an average 1.068×1.068$\times$1.068 × code size reduction on cBench with a single trial after training, surpassing random search methods that require significantly more time—even when accounting for CompilerDream’s training time. This highlights the effectiveness of CompilerDream’s model-based RL training process in finding superior optimization passes, making it a promising component in building a general code optimizer. Moreover, when combined with guided search, CompilerDream outperforms the top method on the leaderboard while using less wall time.
4.3 General Value Prediction
We then demonstrate that the large-scale trained world model in CompilerDream can generalize to accurately simulate the optimization process of unseen input programs, thereby supporting the training of a general optimization agent. To illustrate this, we utilize our trained world model to estimate the outcome of applying an optimization sequence, a task named value prediction.
Implementation
We adopt the state-of-the-art Coreset-NVP method (Liang et al. 2023), which evaluates pass sequences within a fixed candidate set (the coreset with 625 sequence prefixes) and selects the best prefix via compiler validation. Coreset-NVP employs a neural network to directly predict the normalized value of sequences, which is learned via supervised learning. Our Coreset-CompilerDream method leverages a world model trained on CodeContests, to simulate the optimization process and accumulate predicted rewards as the value. Both our method and the baseline method Coreset-NVP are tested on four datasets orthogonal to their training sets. The number of compiler validation calls is set the same across methods to ensure a fair comparison. Additional implementation details are available in Appendix B.8.
Method Dataset
cBench CHStone MiBench NPB
Coreset-NVP 1.028 1.101 1.003 1.085
Coreset
-CompilerDream
1.038 1.101 1.017 1.140
Coreset-Oracle 1.041 1.106 1.020 1.159
Table 2: General value prediction results: The geometric mean of code size reduction achieved by the best sequences selected by different methods across 4 datasets. Results of Coreset-NVP and Coreset-Oracle is reported from Coreset-NVP’s original paper (Liang et al. 2023).
Results
Table 2 shows the geometric mean reduction achieved by different methods. Despite the test datasets being distinct from the training set, CompilerDream outperforms Coreset-NVP across three datasets, highlighting its strong compiler modeling and generalization capabilities. On MiBench, CompilerDream performs nearly optimally compared to the Oracle baseline, which uses a brute-force search and serves as the problem’s upper bound. Both our method and Coreset-NVP achieve an average reduction of $1.101$ on CHStone, close to the upper bound of $1.06$.
4.4 General Code Optimization
We finally aim to demonstrate that CompilerDream can learn a general optimization agent, capable of end-to-end generating effective optimization sequences for a wide variety of programs unseen during training. This scenario represents a typical daily use case where the optimization algorithm has limited time to compute the best solution for each of the diverse programs. Consequently, all methods generate only one optimization sequence per program in the test datasets unless otherwise specified.
Figure 4: Histograms of large-scale evaluations comparing CompilerDream and PPO on the FormAI dataset.
Figure 6: Analysis. Evaluations of different RL algorithms: (a) Learning curves of various RL algorithms, measured by the geometric mean of code size reduction on the CodeContests validation set. A Gaussian filter ($\sigma=2.0$) is applied to enhance the visualization of trends. (b) Zero-shot generalization capabilities of different RL algorithms on various test benchmark datasets. Effect of training dataset: (c) Test performance of CompilerDream trained on CodeContests and Csmith. Bars indicate the standard deviation.
Implementation
We compare CompilerDream with the following baselines: a random pass sequence, an autotuning approach using random search, LLVM’s -O0 and -Oz flag, and a state-of-the-art learning-based method. The random search conducts hundreds of trials within a time budget similar to CompilerDream’s. The -Oz flag represents LLVM’s highest level of code size optimization while the -O0 flag represents no optimization. The learning-based method using Proximal Policy Optimization (PPO) (Schulman et al. 2017) and CompilerDream are trained on the large-scale CodeContests dataset and zero-shot generalized to unseen test programs. Following the state-of-the-art approach, all methods operate within a limited action space.
Results
Figure 3 shows CompilerDream’s performance in reducing code size, measured by the geometric mean of IR instruction count reduction. Without specific in-domain training, CompilerDream outperforms -Oz in all but two test benchmarks in a single trial and surpasses PPO in all but the BLAS datasets. It also exceeds random search within a similar time budget on most datasets, except Linux. The minimal performance differences on the BLAS and Linux datasets suggest they are already highly optimized.
Moreover, CompilerDream’s zero-shot generalization matches or exceeds in-domain training. This advantage is especially notable in the NPB dataset, where data sparsity challenges in-domain agents, yet CompilerDream achieves an additional $3\%$ code size reduction. CompilerDream also generalizes well to new languages, evidenced in its performances on the BLAS and NPB datasets of Fortran programs. Further, on the AI-generated Objective-C dataset, CompilerDream achieves an average code size reduction of 1.027$\times$, reaching up to 2.87$\times$ in certain instances.
On the large-scale FormAI test set (Figure 4), CompilerDream outperforms PPO, matching or outperforming -Oz on more programs and achieving higher optimization levels, demonstrating its superior and consistent performance.
4.5 Analysis
Comparison with model-free methods
We further evaluate the sample efficiency and zero-shot generalization abilities of CompilerDream based on the world model algorithm, Dreamer, against a range of model-free counterparts, including PPO (Schulman et al. 2017), DQN (Mnih et al. 2015), A2C (Mnih et al. 2016), APEX (Horgan et al. 2018), and IMPALA (Espeholt et al. 2018; Schulman et al. 2017). Figure 6a illustrates that while PPO is the strongest model-free baseline, CompilerDream outperforms it, learning an order of magnitude faster with fewer compiler interactions. As Figure 6b demonstrates, CompilerDream also excels in generalizing to unseen benchmarks, supporting our hypothesis that world model-based agents better capture the environment’s structure and enhance zero-shot generalization.
Effect of training dataset
To assess the impact of the CodeContests dataset on the generalization ability, we compared it with the commonly used Csmith (Yang et al. 2011) dataset, a large LLVM IR dataset generated by rules. We train CompilerDream on both datasets and the results shown in Figure 6 indicate that CompilerDream trained on CodeContests significantly outperforms the Csmith-trained version across all five test datasets. This advantage is particularly evident on the manually curated cBench (Fursin et al. 2007) and CHStone (Hara et al. 2008) datasets, demonstrating that the CodeContests-trained model can generalize more effectively to human-written programs.
Program showcase
5 Discussion
We aim to tackle the major challenge of zero-shot generalization that arises when applying learning-based code optimization. We thus introduce the CompilerDream approach that leverages the sample efficiency and generalization capabilities of world model-based reinforcement learning, as well as the strengths of a diverse, high-quality training program dataset. Our results have shown that CompilerDream, with its powerful generalization ability, attains optimization performance that outperforms existing methods, in multiple problem scenarios under the challenging area of phase ordering. There is substantial scope for further exploration, including expansion of the training dataset, scaling up the compiler world model, optimizing multiple objectives like execution time, and enriching feature and action spaces with deeper expert knowledge or large language models.
Appendix A Extended Related Work
Here we provide an overview of several topics in compiler optimization that our work did not focus on but are nonetheless important:
Feature Extraction